OpenStudy (anonymous):

am i correct? 10x^6 + 8x^4 / 2x^2 = 9x^12

jimthompson5910 (jim_thompson5910):

I'm assuming the original expression is \[\large \frac{10x^6 + 8x^4}{2x^2}\] right?

OpenStudy (anonymous):

yes jim

jimthompson5910 (jim_thompson5910):

so you can either break up the fraction and reduce, or you can factor out 2x^2 from the numerator and reduce

OpenStudy (agent0smith):

Break up the fraction like so:\[\Large \frac{10x^6 + 8x^4}{2x^2} = \frac{10x^6}{2x^2}+ \frac{8x^4}{2x^2}\]

jimthompson5910 (jim_thompson5910):

if you were to factor out 2x^2, you would get \[\large \frac{10x^6 + 8x^4}{2x^2}\] \[\large \frac{2x^2(5x^4 + 4x^2)}{2x^2}\] hopefully you see what to do next

OpenStudy (anonymous):

got it thx

jimthompson5910 (jim_thompson5910):

you're welcome